OpenStudy (crashonce):

In a GP the sum of the first two terms is 70 and the sum of the first six terms is 910. Find the sum of the first four terms

OpenStudy (crashonce):

@jim_thompson5910 an easy one

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (crashonce):

i cant get it

jimthompson5910 (jim_thompson5910):

I'm still thinking it through

OpenStudy (shinalcantara):

i got 280 but i still have to figure it out.. I'm using shift solve though..

jimthompson5910 (jim_thompson5910):

well if you use \[\Large S_{n} = a*\frac{1-r^n}{1-r}\] you can set up \[\Large S_{2} = a*\frac{1-r^2}{1-r} = 70\] \[\Large S_{6} = a*\frac{1-r^6}{1-r} = 910\]

OpenStudy (crashonce):

ive done that but my anwer is wierd

jimthompson5910 (jim_thompson5910):

Divide the two equations to get \[\Large \frac{1-r^2}{1-r^6} = \frac{70}{910}\] now solve for r

OpenStudy (shinalcantara):

\[S_n = a_1\frac{ 1-r^n }{ 1-r }\] \[70 = a_1\frac{ 1-r^2 }{ 1-r }\] \[a_1 = \frac{ 70(1-r) }{ 1-r^2 }\] \[910 = a_1\frac{ 1-r^6 }{ 1-r }\] \[910 = \left( \frac{ 70(1-r) }{ 1-r^2 } \right)\left( \frac{ 1-r^6 }{ 1-r } \right)\] \[\frac{ 910 }{ 70 } = \frac{ 1-r^6 }{ 1-r^2 }\] \[13 = \frac{ 1-r^6 }{ 1-r^2 }\] is there a way to get the value of 'r' without the help of a shift solve function in the calculator @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

it turns out that \[\Large 1-r^6 = 1-(r^3)^2\] \[\Large 1-r^6 = (1-r^3)(1+r^3)\] \[\Large 1-r^6 = (1-r)(1+r+r^2)(1+r^3)\] \[\Large 1-r^6 = (1-r)(1+r+r^2)(1+r)(1-r+r^2)\] \[\Large 1-r^6 = (1-r)(1+r)(1+r+r^2)(1-r+r^2)\] \[\Large 1-r^6 = (1-r)(1+r)(r^4+r^2+1)\] so \[\Large \frac{1-r^2}{1-r^6} = \frac{70}{910}\] \[\Large \frac{(1-r)(1+r)}{(1-r)(1+r)(r^4+r^2+1)} = \frac{70}{910}\] \[\Large \frac{1}{r^4+r^2+1} = \frac{70}{910}\] \[\Large \frac{1}{r^4+r^2+1} = \frac{1}{13}\] \[\Large r^4+r^2+1 = 13\] This is a biquadratic which can be solved exactly

OpenStudy (crashonce):

how do you solve it, never seen this b4

jimthompson5910 (jim_thompson5910):

let z = r^2, so z^2 = r^4 you'll get \[\Large z^2 + z+1 = 13\] solve for z, then use the solutions for z to get the solutions for r

OpenStudy (crashonce):

how don't get it isnt it no answer

jimthompson5910 (jim_thompson5910):

what solutions for z did you get?

jimthompson5910 (jim_thompson5910):

use the quadratic formula